1. For what values of h the vectors


⟶ [1 ⟶ [-1 ⟶ [5

u1 = -3 u2 = 9 u3 = -7

-2] -6] h]


are linearly independent? (Show all working)


1
Expert's answer
2022-03-19T02:39:43-0400

Solution


Given the three vectors are


"\vec{u_{1}}=[1, -3, -2]"


"\vec{u_{2}}=[5, 9, -6]"


"\vec{u_{3}}=[5, -7, h]"


Now the three vectors will be linearly independent when for three scalars, "a, b" and "c" , we can write,


"a\vec{u_{1}}+b\vec{u_{2}}+c\vec{u_{3}}=0" if and only iff when "a=b=c=0"


Therefore,


"a\vec{u_{1}}+b\vec{u_{2}}+c\vec{u_{3}}=a[1, -3, -2]+b[5, 9, -6]+c[5, -7, h]=0"


"a[1, -3, -2]+b[5, 9, -6]+c[5, -7, h]=[0, 0, 0]"


"[a, -3a, -2a]+[5b, 9b, -6b]+[5c, -7c, hc]=[0, 0, 0]"


"a+5b+5c=0\\
-3a+9b-7c=0\\
-2a-6b+hc=0\\"


Solving these equations,





since for the linearly independent, "a=b=c=0" , which requires,


"h+\frac{26}{3}\neq0"


"h\neq\frac{26}{3}"


Hence other than this value, the three vectors will be linearly independent.